# Deep learning-derived arterial input function

### v1

@@ 5.3 Qualitative and quantitative analysis
Our study contrasts the proposed DLIF framework with the established traditional EPICA method [40, 41]. This method relies on decomposing uptake values categorized into different regions through independent component analysis. Originally, EPICA required invasive arterial sampling to obtain the peak value for scaling the estimated IDIF. However, for the purposes of our evaluation, we use the peak value from the actual ground-truth AIF as the scaling factor. Furthermore, given that DLIF represents a novel approach by integrating deep learning for input function estimation, it stands without direct predecessors employing a similar methodology. As a result, we benchmark the DLIF framework against two prominent DNNs in image classification, ResNet50 [23] and ConvNeXt [34]. These models are adapted to produce discrete input function values at specified time points. They serve as strong comparatives to the ViT architecture within the DLIF framework, providing a comprehensive comparison within the deep learning domain.
@@
Fig. 6 shows the qualitative results of the DLIFs under the three top performing configurations, "Exp.", "Gaussian+Peak", and "Gaussian+Peak+Sparse", alongside the input functions estimated by EPICA, ResNet50, and ConvNeXt for four different subjects. The first two subjects are characterized by HAB genotypes, while the last two are MAB genotypes. The EPICA method noticeably falls short, failing to closely match the true AIF shapes. This is particularly evident in the zoom-in views, where EPICA often inaccurately identifies peak time points, resulting in jagged and inferior amplitude alignment in the input functions’ tails. This limitation is likely due to its reliance on an unsupervised strategy that estimates AIFs by decomposing voxel values across the dynamic PET scan sequence using independent component analysis, which may be affected by noise presented in the PET images. In contrast, the DNN-based supervised learning approaches show significantly better adherence to the ground truth AIF shapes. Within this group, ResNet50 exhibits the least effectiveness in capturing both the shape and amplitude of the AIFs. ConvNeXt, while adept at pinpointing peak values and time points, tends to produce results that are not as smooth, with less accurate amplitude matching post-peak. Upon closer examination, highlighted in the second and third rows of Fig. 6, it is evident that the proposed DLIF framework generates AIFs with a smoother consistency and a closer match to the ground truth across both patient genotypes compared to other methods, which show greater fluctuations in their estimated input functions.
@@
Fig. 7: This figure illustrates the parametric maps (left) reconstructed based on AIFs and their percentage differences relative to the ground truth (right). Each row represents a different subject, and each column represents a different AIF estimation method. The subjects are grouped according to their genotypes: C/C or C/T.
@@
Table 1 showcases the quantitative analysis comparing various methods, with the top performances emphasized in bold. Statistical significance tests, specifically the Wilcoxon signed-rank tests, were conducted between the top-performing non-DLIF model, ConvNeXt, and the top DLIF models to assess significance, with Bonferroni correction applied to results with the identical scores. The quantitative results for EPICA, which correlate with its qualitative evaluations, indicate it as the least effective method. It shows mean $\boldsymbol{r}$ values of approximately 0.5 and suboptimal mean IoU values of about 0.35, demonstrating poor correspondence in both shape and integral matching of the input functions. In contrast, the deep learning-based methods surpass EPICA across all metrics, with the proposed DLIF models showing exceptional performance, particularly achieving mean $\boldsymbol{r}$ values of 0.94 for both HAB and MAB patients, which indicates strong shape agreement with the ground truth AIFs from arterial blood sampling. Look at the other metrics, the "Gaussian+Peak" configuration of the proposed DLIF demonstrated statistically significant improvements in IoU and RMSE for the tail of the AIFs (i.e., IoU (30+)), with mean IoU values of 0.76 and 0.77, and mean RMSE values of 0.06 and 0.07, respectively, for HAB and MAB genotypes. The "Gaussian+Peak+Sparse" configuration yielded the best peak matching, showing the least peak bias with values of 0.31 and 0.26 for HAB and MAB genotypes respectively. In the peak regions (i.e., <30 mins or 30-), all deep learning-based methods performed reasonably well, with the proposed DLIFs achieving the highest mean IoU and RMSE values, surpassing ResNet50 significantly and being slight better than the results of ConvNeXt.
@@
Fig. 8: Visualization of attention maps provided by the ViT of the proposed DLIF framework, showcasing nine distinct configurations. These range from direct estimation of the AIFs ("Direct Est.") to implementations using basis functions, both with and without the integration of peak estimation and sparsity constraints.
@@ 5.4 Logan graphical analysis
We also undertook secondary, indirect comparisons using the parametric maps generated from the estimated input functions. These brain distribution volume maps ($V_{T}$) were derived using Logan graphical analysis, as detailed in the Methods section [35]. Figure 7 provides qualitative comparisons of these parametric maps for eight patients spanning two genotypes: the upper four rows illustrate HAB genotypes, and the lower four rows show MAB genotypes. The left panel displays the parametric maps, the middle panel highlights absolute differences between estimated and ground truth values, and the right panel features scatter plots comparing voxel-wise estimations to ground truth values.
@@
  EPICA 3.51$\pm$6.32 3.65$\pm$6.55
ResNet50 1.24$\pm$1.00 1.27$\pm$1.02
ConvNeXt 0.67$\pm$0.64 0.69$\pm$0.66
DLIF-Exp. 0.75$\pm$0.73 0.77$\pm$0.75
DLIF-Gaussian+Peak 0.66$\pm$0.69 0.68$\pm$0.70
DLIF-Gaussian+Peak+Sparse 0.71$\pm$0.80 0.73$\pm$0.81
@@ 6 Discussion
To further compare the estimated and ground truth AIFs, we introduced a novel evaluation metric based on the ratio of the integrated area of overlap, or IoU. This metric emphasizes the integral areas under the curves, which is physiologically important for determining total distribution volume, rather than focusing solely on matching peak values at specific time points. Based on the $\boldsymbol{r}$ and IoU metrics, we identified three particularly effective configurations of the DLIF on the validation dataset: "Exp.", "Gaussian+Peak", and "Gaussian+Peak+Sparse". We compared these configurations against various baseline methods on the test set. Qualitative results in Fig. 6 confirmed the capability of DLIF to match the peak of ground truth while producing a smooth tail that closely matches the original data. Other methods resulted in a less smooth tail, with the traditional method EPICA performing the lowest, as its estimated input function struggled to match the ground truth even though it required the ground truth peak value as a scaling factor. Quantitative results shown in Table 1 further validate our findings, with DLIF configurations significantly outperforming others. When comparing the parametric maps estimated using the AIFs predicted by various methods, all DLIF configurations showed a good match with the ground truth, as indicated by the absolute difference images and the scatter plots shown in Fig. 7. The "Gaussian+Peak" configuration demonstrated the least quantitative error, as detailed in Table 2.
@@
DL models are often described as "black-box" models due to their limited ability to explain how specific estimates are derived. However, a notable advantage of using ViT within the DLIF framework is its inherent capability to generate attention maps, elucidating the regions of focus during prediction. Figure 8 illustrates attention map visualizations from the final self-attention layer of the ViT for sixteen randomly selected datasets across various DLIF configurations. Given that DLIF aims to directly estimate metabolite-corrected AIFs from dynamic PET images reflecting radiotracer uptake in the brain, it is expected that the ViT predominantly attends to regions of high uptake (higher intensity voxels). In the "Direct Est." configuration, while the ViT’s attention is generally centered on the brain tissue, it occasionally diverts to less pertinent background regions. The introduction of basis functions without peak estimation or sparsity enforcement, seen in the "Gaussian" and "Exp." configurations, led to attention mechanisms occasionally fixating on confined background regions, making interpretation non-intuitive. Imposing sparsity constraints slightly refined the attention mechanisms’ focus within the brain, although some attention still lingered on the background. In contrast, integrating peak estimation markedly improved focus on brain regions, with attention spreading to areas of high uptake, especially evident in the "Gaussian+Peak" and "Exp.+Peak" configurations. However, adding a sparsity constraint redirected the attention mechanisms towards the background again, possibly because enforcing sparsity prompted the ViT to leverage small values in these regions to assign smaller weights to certain basis functions.
Notably, although this study focused specifically on TSPO-PET imaging of the brain, the DLIF framework’s image-based approach is highly adaptable. It can be readily extended to other anatomical sites and radiotracers by retraining or fine-tuning the DNN. Additionally, while the presented data were acquired from a single PET scanner, employing similar retraining and calibration procedures enables compatibility with data acquired on other PET scanners and imaging protocols. In summary, our results underscore the effectiveness of supervised deep-learning methodologies, such as DLIF, in providing accurate, fully non-invasive estimates of metabolite-corrected AIFs.
@@ 7 Conclusion
In this study, we introduced DLIF, a deep learning-based framework designed to non-invasively estimate AIF, entirely eliminating the necessity of invasive arterial blood sampling. The DLIF framework incorporates prior physiological knowledge from existing literature by representing AIFs using continuous basis functions, such as Gaussian and exponential-sigmoid functions. By leveraging the flexibility of DNNs, DLIF accurately models complex temporal dynamics through an overcomplete representation of these basis functions, producing input functions as combinations of individually predicted parameters. Unlike traditional methods, including IDIF and PBIF, which yield a non-personalized estimation of AIF [66, 2] or often require a scaling factor from single or multiple time points invasive arterial or venous blood sampling to calculate AIFs for patients [65, 56, 40, 41, 42, 31, 7], DLIF completely avoids such invasive procedures. It offers fast and robust estimation of AIFs using only the subjects’ dynamic PET sequences. This makes the DLIF framework a promising tool for broader adoption of dynamic PET imaging in routine clinical settings, replacing current semi-quantitative analysis methods. This advancement could substantially enhance diagnostic accuracy for neurological conditions such as Alzheimer’s disease and Parkinson’s disease, particularly during early-stage evaluations.

### v2

@@ 5.3 Qualitative and quantitative analysis
Our study contrasts the proposed DLIF framework with the established traditional EPICA method [naganawa2005extraction, naganawa2005omission]. This method relies on decomposing uptake values categorized into different regions through independent component analysis. Originally, EPICA required invasive arterial sampling to obtain the peak value for scaling the estimated IDIF. However, for the purposes of our evaluation, we use the peak value from the actual ground-truth AIF as the scaling factor. Furthermore, given that DLIF represents a novel approach by integrating deep learning for input function estimation, it stands without direct predecessors employing a similar methodology. As a result, we benchmark the DLIF framework against two prominent DNNs in image classification, ResNet50 [he2016deep] and ConvNeXt [liu2022convnet]. These models are adapted to produce discrete input function values at specified time points. They serve as strong comparatives to the ViT architecture within the DLIF framework, providing a comprehensive comparison within the deep learning domain.
@@
Fig. 6 shows the qualitative results of the DLIFs under the three top performing configurations, “Exp.”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse”, alongside the input functions estimated by EPICA, ResNet50, and ConvNeXt for four different subjects. The first two subjects are characterized by HAB genotypes, while the last two are MAB genotypes. The EPICA method noticeably falls short, failing to closely match the true AIF shapes. This is particularly evident in the zoom-in views, where EPICA often inaccurately identifies peak time points, resulting in jagged and inferior amplitude alignment in the input functions’ tails. This limitation is likely due to its reliance on an unsupervised strategy that estimates AIFs by decomposing voxel values across the dynamic PET scan sequence using independent component analysis, which may be affected by noise presented in the PET images. In contrast, the DNN-based supervised learning approaches show significantly better adherence to the ground truth AIF shapes. Within this group, ResNet50 exhibits the least effectiveness in capturing both the shape and amplitude of the AIFs. ConvNeXt, while adept at pinpointing peak values and time points, tends to produce results that are not as smooth, with less accurate amplitude matching post-peak. Upon closer examination, highlighted in the second and third rows of Fig. 6, it is evident that the proposed DLIF framework generates AIFs with a smoother consistency and a closer match to the ground truth across both patient genotypes compared to other methods, which show greater fluctuations in their estimated input functions.
@@
Fig. 7: This figure illustrates the parametric maps (left) reconstructed based on AIFs and their percentage differences relative to the ground truth (right). Each row represents a different subject, and each column represents a different AIF estimation method. The subjects are grouped according to their genotypes: C/C or C/T. For visualization, subject-specific color scales are used to account for the large inter-subject variability in intensity ranges.
@@
Table 1 showcases the quantitative analysis comparing various methods, with the top performances emphasized in bold. Statistical significance tests, specifically the Wilcoxon signed-rank tests, were conducted between the top-performing non-DLIF model, ConvNeXt, and the top DLIF models to assess significance, with Bonferroni correction applied to results with the identical scores. The quantitative results for EPICA, which correlate with its qualitative evaluations, indicate it as the least effective method. It shows mean $\boldsymbol{r}$ values of approximately 0.5 and suboptimal mean IoU values of about 0.35, demonstrating poor correspondence in both shape and integral matching of the input functions. In contrast, the deep learning-based methods surpass EPICA across all metrics, with the proposed DLIF models showing exceptional performance, particularly achieving mean $\boldsymbol{r}$ values of 0.94 for both HAB and MAB patients, which indicates strong shape agreement with the ground truth AIFs from arterial blood sampling. Look at the other metrics, the “Gaussian+Peak” configuration of the proposed DLIF demonstrated statistically significant improvements in IoU and RMSE for the tail of the AIFs (i.e., IoU (30+)), with mean IoU values of 0.76 and 0.77, and mean RMSE values of 0.06 and 0.07, respectively, for HAB and MAB genotypes. The “Gaussian+Peak+Sparse” configuration yielded the best peak matching, showing the least peak bias with values of 0.31 and 0.26 for HAB and MAB genotypes respectively. In the peak regions (i.e., <30 mins or 30-), all deep learning-based methods performed reasonably well, with the proposed DLIFs achieving the highest mean IoU and RMSE values, surpassing ResNet50 significantly and being slight better than the results of ConvNeXt.
@@
Fig. 8: Visualization of attention maps provided by the ViT of the proposed DLIF framework, showcasing nine distinct configurations. These range from direct estimation of the AIFs (“Direct Est.”) to implementations using basis functions, both with and without the integration of peak estimation and sparsity constraints.
@@ 5.4 Logan graphical analysis
We also undertook secondary, indirect comparisons using the parametric maps generated from the estimated input functions. These brain distribution volume maps ($V_{T}$) were derived using Logan graphical analysis, as detailed in the Methods section [logan1990graphical]. Figure 7 provides qualitative comparisons of these parametric maps for eight patients spanning two genotypes: the upper four rows illustrate HAB genotypes, and the lower four rows show MAB genotypes. The left panel displays the parametric maps, the middle panel highlights absolute differences between estimated and ground truth values, and the right panel features scatter plots comparing voxel-wise estimations to ground truth values.
@@
[…]
@@ 6.1 Performance of DLIF Configurations and Evaluation Metrics
To further compare the estimated and ground truth AIFs, we introduced a novel evaluation metric based on the ratio of the integrated area of overlap, or IoU. This metric emphasizes the integral areas under the curves, which is physiologically important for determining total distribution volume, rather than focusing solely on matching peak values at specific time points. Based on the $\boldsymbol{r}$ and IoU metrics, we identified three particularly effective configurations of the DLIF on the validation dataset: “Exp.”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse”. We compared these configurations against various baseline methods on the test set. Qualitative results in Fig. 6 confirmed the capability of DLIF to match the peak of ground truth while producing a smooth tail that closely matches the original data. Other methods resulted in a less smooth tail, with the traditional method EPICA performing the lowest, as its estimated input function struggled to match the ground truth even though it required the ground truth peak value as a scaling factor. Quantitative results shown in Table 1 further validate our findings, with DLIF configurations significantly outperforming others. When comparing the parametric maps estimated using the AIFs predicted by various methods, all DLIF configurations showed a good match with the ground truth, as indicated by the absolute difference images and the scatter plots shown in Fig. 7. The “Gaussian+Peak” configuration demonstrated the least quantitative error, as detailed in Table 2.
6.2 Genotype-Specific Analysis
@@ 6.3 Architectural Flexibility of DLIF
Table 4: Quantitative results of different methods from the generalizability and transfer learning studies. The results are based on five-fold cross-validation. Note that the time intervals “30-” and “30+” refer to the periods of 0-30 minutes and 60-90 minutes, respectively. The top-performing scores are highlighted in bold. The symbol “*” denotes a statistically significant improvement between without and with transfer learning (i.e., fine-tuning), with a $p$-value < 0.05, as determined by a Wilcoxon signed-rank test.
                 [11C]DPA-713 PET data acquired on the Siemens Biograph mCT scanner (New Scanner)
Without fine-tuning (zero-shot) With fine-tuning
Metrics DLIF-Exp. DLIF-Gaussian+Peak DLIF-Gaussian+Peak+Sparse DLIF-Exp. DLIF-Gaussian+Peak DLIF-Gaussian+Peak+Sparse
[…]
[11C]CPPC PET data acquired on the Siemens Biograph mCT scanner (New tracer & scanner)
Without fine-tuning (zero-shot) With fine-tuning
Metrics DLIF-Exp. DLIF-Gaussian+Peak DLIF-Gaussian+Peak+Sparse DLIF-Exp. DLIF-Gaussian+Peak DLIF-Gaussian+Peak+Sparse
[…]
6.4 Interpretability through Attention Maps
@@
DL models are often described as “black-box” models due to their limited ability to explain how specific estimates are derived. However, a notable advantage of using ViT within the DLIF framework is its inherent capability to generate attention maps, elucidating the regions of focus during prediction. Figure 8 illustrates attention map visualizations from the final self-attention layer of the ViT for sixteen randomly selected datasets across various DLIF configurations. Given that DLIF aims to directly estimate metabolite-corrected AIFs from dynamic PET images reflecting radiotracer uptake in the brain, it is expected that the ViT predominantly attends to regions of high uptake (higher intensity voxels). In the “Direct Est.” configuration, while the ViT’s attention is generally centered on the brain tissue, it occasionally diverts to less pertinent background regions. The introduction of basis functions without peak estimation or sparsity enforcement, seen in the “Gaussian” and “Exp.” configurations, led to attention mechanisms occasionally fixating on confined background regions, making interpretation non-intuitive. Imposing sparsity constraints slightly refined the attention mechanisms’ focus within the brain, although some attention still lingered on the background. In contrast, integrating peak estimation markedly improved focus on brain regions, with attention spreading to areas of high uptake, especially evident in the “Gaussian+Peak” and “Exp.+Peak” configurations. However, adding a sparsity constraint redirected the attention mechanisms towards the background again, possibly because enforcing sparsity prompted the ViT to leverage small values in these regions to assign smaller weights to certain basis functions.
6.5 Generalizability of the DLIF Framework
We evaluated the generalizability of the proposed DLIF framework by testing it on two new tasks designed to probe its adaptability and potential for transfer learning. First, to assess robustness to hardware variability, we applied DLIF to 21 [11C]DPA-713-TSPO-PET dynamic scans acquired on a newer generation scanner (Siemens Biograph mCT, a time-of-flight whole body system). Second, to examine adaptability across tracers, we applied DLIF to 10 [11C]CPPC-CSF1R-PET dynamic scans acquired with [11C]CPPC on the Siemens Biograph mCT system [coughlin2022first, horti2019pet, mills2025exploring]. [11C]CPPC binds to a different target, CSF1R, and exhibits pharmacokinetic properties very distinct from [11C]DPA-713, which binds to TSPO. Compared to [11C]DPA-713, [11C]CPPC exhibits a slower washout and lower peak in the AIF, representing behavior not encountered during the original DLIF training. All PET images were normalized to SUVs and the datasets included associated metabolite-corrected AIFs.
Fig. 9: Qualitative results for the generalizability and transfer learning tasks. The left panel shows AIFs and DLIF estimates from [11C]DPA-713-TSPO-PET data acquired on a newer-generation scanner (Siemens Biograph mCT), while the right panel presents results from [11C]CPPC-CSF1R-PET data acquired on a scanner of the same model type. Each column corresponds to a unique subject. The first row presents the full estimated input functions, and the second and third rows show zoomed-in views of the 0–7 minute and 22–90 minute intervals, respectively. “w/o FT” and “w/ FT” in the legend indicate results obtained without and with fine-tuning, respectively.
For both tasks, we first evaluated the zero-shot capability of DLIF framework by directly applying the model trained on [11C]DPA-713 data from HRRT scanners. We then fine-tuned the model by allowing only the estimation heads and the final Transformer block to remain trainable, while freezing all other parameters. The model was fine-tuned for 100 epochs with a learning rate of 1e-4. This transfer learning strategy reduces the number of trainable parameters, enabling effective adaptation to new domains under limited data availability while mitigating overfitting. We performed five-fold cross-validation for both tasks, with each fold using 8:2 split for training and evaluation.
6.5.1 New scanner
The left panel of Fig. 9 shows two examples of DLIF estimates compared with the ground-truth AIFs. Despite the change in scanners, the DLIF outputs without fine-tuning closely matched the reference AIFs, though with slightly lower peak values. This observation is corroborated by the quantitative results in the upper panel of Table 4, where zero-shot application of the DLIF framework to the new scanner data achieved performance comparable to that reported in Table 1, which was obtained from test images drawn from the same distribution as the training dataset. Specifically, the correlation coefficient ($\boldsymbol{r}$) remained around 0.95, IoU values exceeded 0.7, and both RMSE and peak bias were relatively small across all three DLIF variants. These findings indicate that the proposed DLIF framework achieves strong zero-shot performance for [11C]DPA-713 scans acquired on different scanners, suggesting that it captures prior knowledge of [11C]DPA-713 AIFs in a manner largely invariant to scanner differences. Further investigation is needed to determine whether this robustness extends to additional scanner types and acquisition protocols.
Although the zero-shot performance was already strong, additional improvements were achieved through fine-tuning. Qualitative results in Fig. 9 demonstrate a closer match between the DLIF estimates and the ground-truth AIFs after fine-tuning. Quantitative scores in Table 4 also showed statistically significant gains, particularly in shape and peak matching, with the best performance achieved by the “Gaussian+Peak+Sparse” variant. Improvements were most pronounced in peak-related metrics: IoU and RMSE within the first 30 minutes, as well as peak bias, all showed significant improvement relative to their zero-shot counterparts, whereas performance during the later 60–90 minute period remained comparable.
6.5.2 New tracer & scanner
The right panel of Fig. 9 illustrates results for [11C]CPPC. In this case, the pretrained DLIF model failed to produce accurate predictions, with both the peak amplitude and overall AIF shape deviating substantially from the ground truth. Quantitative results in the lower panel of Table 4 confirm this observation: without fine-tuning, the correlation coefficient ($\boldsymbol{r}$) remained in the suboptimal range of 0.7–0.8, IoU values were consistently below 0.7, and both RMSE and peak bias were relatively high. These findings support our hypothesis that a model trained exclusively on [11C]DPA-713 data develops prior knowledge specific to [11C]DPA-713 AIF characteristics, which enables robustness to scanner variability but limits generalization to tracers unseen from training with distinct pharmacokinetics such as [11C]CPPC.
The key question is whether the DLIF framework trained on [11C]DPA-713 can be effectively transferred to [11C]CPPC, which was also acquired on a different scanner (mCT instead of HRRT). The results demonstrate that it can: with only eight subjects used for fine-tuning in each fold, performance improved substantially. This improvement is evident in both the qualitative results (Fig. 9) and the quantitative results (Table 4), where the scores became comparable to those obtained on [11C]DPA-713 data (Table 1). Consistently, most metrics also showed statistically significant improvements relative to their zero-shot counterparts.
In this work, the metabolite-corrected AIFs were used as ground truth for training; therefore, the DLIF framework directly predicts metabolite-corrected AIFs without explicitly modeling metabolite kinetics. For tracers that do not produce metabolites, require whole-blood rather than metabolite-corrected input functions, or exhibit tracer-specific metabolite behavior such as metabolites crossing the blood–brain barrier, additional training or transfer learning using the appropriate input functions would be required. Despite these tracer-specific considerations, the underlying DLIF framework is general and can be adapted beyond the current scope. Although this study focused on TSPO-PET imaging of the brain, the framework can, with fine-tuning on small tracer-specific datasets or retraining with new data, be extended to other tracers, anatomical regions, and imaging protocols. Overall, these findings highlight the potential of DLIF to provide accurate and fully non-invasive estimates of arterial input functions across diverse PET imaging applications.
6.6 Compartment Analysis using DLIF
Given the close agreement between the DLIF- and AIF-derived curves, we further investigated whether DLIF could be extended beyond Logan graphical analysis to support full compartmental modeling. As a proof of concept, we analyzed a representative [11C]DPA-713 scan acquired on the Siemens Biograph mCT scanner (as described in Section 6.5.1) using the fine-tuned DLIF model for AIF estimation. A two-tissue compartment model was applied, and the total distribution volume ($V_{T}$) values across multiple brain regions were compared with those obtained using ground-truth AIFs. The results are presented in B (Fig. 10), which shows a correlation plot where the $x$-axis represents $V_{T}$ estimates derived from ground-truth AIFs and the $y$-axis represents those derived from DLIF, with each circle corresponding to a distinct brain region. The correlation between DLIF- and AIF-based $V_{T}$ estimates was close to unity (dashed line), indicating strong agreement. An example fit for the frontal cortex further demonstrates that the DLIF- and AIF-driven model fits closely overlap. This proof-of-concept analysis illustrates that DLIF-derived AIFs can reliably support compartmental kinetic modeling, highlighting their potential utility beyond simplified approaches such as Logan graphical analysis.
@@ 7 Conclusion
In this study, we introduced DLIF, a deep learning-based framework designed to non-invasively estimate AIF, entirely eliminating the necessity of invasive arterial blood sampling. The DLIF framework incorporates prior physiological knowledge from existing literature by representing AIFs using continuous basis functions, such as Gaussian and exponential-sigmoid functions. By leveraging the flexibility of DNNs, DLIF accurately models complex temporal dynamics through an overcomplete representation of these basis functions, producing input functions as combinations of individually predicted parameters. Unlike traditional methods, including IDIF and PBIF, which yield a non-personalized estimation of AIF [Zanotti_Fregonara_2012, Boutin_2007] or often require a scaling factor from single or multiple time points invasive arterial or venous blood sampling to calculate AIFs for patients [zanotti2011image, Takikawa_1993, naganawa2005extraction, naganawa2005omission, naganawa2008robust, litton1997input, chen1998noninvasive], DLIF completely avoids such invasive procedures. It offers fast and robust estimation of AIFs using only the subjects’ dynamic PET sequences. This makes the DLIF framework a promising tool for broader adoption of dynamic PET imaging in routine clinical settings, replacing current semi-quantitative analysis methods. This advancement could substantially enhance diagnostic accuracy for neurological conditions such as Alzheimer’s disease and Parkinson’s disease, particularly during early-stage evaluations.